for the given functions f and g, find the indecated compastion. f(x) = 5/x+8,g(x)= 1/7x (fog)(x)
Answer by stanbon(75887)

find the indecated composition. f(x) = 5/(x+8) , g(x)= (1/7)x
-----------
(fog)(x) = f[g(x)]
= f[(1/7)x]
---
= 5/[(x/7)+8]
---
= 5/[(x+56)/8]
---
= 40/(x+56)
=================
Cheers,
Stan H.